# Canonical decompositions of 3-connected graphs

### v2

@@ 2.4. Tools to prove (ii) and (iii)
We say that a ${}^{+}$ $(C,D)$ of a graph $G$ almost interlaces a star $\sigma$ of ${}^{+}$ of $G$ if $(A,B)\leqslant(C,D)$ or $(A,B)\leqslant(D,C)$ for all $(A,B)\in\sigma$.
The notion of ‘almost interlaces’ is more general than the notion of ‘interlaces’ in two ways: on the one hand, we consider ${}^{+}$, and on the other hand we no longer require that $(A,B)$ and its inverse are not in $\sigma$.
@@
Assume Section 1.7. If a tri-separation $(C,D)$ of $G$ almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$, then $S(C,D)$ contains $v$ or an edge incident with $v$.
@@
Similar to Section 2.4, we prove the following (differences are underlined).
@@
2.5. Proof of (ii)
A key step in the proof of (ii) is to show that the tri-star of the apex-decomposition from Section 1.7 is splitting, see Section 2.5 below. Then we finish the proof similarly to the proof of (i). We start preparing to prove Section 2.5.
@@ 2.5. Proof of (ii)
Assume Section 1.7 with crossing tri-separations. If $O$ has type btt, then the tri-star of ${\mathcal{A}}$ is a splitting star of the set of all totally-nested nontrivial tri-separations of $G$.
@@
By Section 1.8, the tri-star of ${\mathcal{A}}$ consists of totally-nested nontrivial tri-separations, and $v$ is adjacent to all three vertices of $O$. It remains to show that the tri-star of ${\mathcal{A}}$ is splitting. Let $x_{1},x_{2},x_{3}$ denote the vertices of $O$ so that the edge $x_{2}x_{3}$ is bold. Let $B_{\ell}$ denote the unique leaf-bag of ${\mathcal{A}}$; so $\{x_{2},x_{3}\}$ is the adhesion set of $B_{\ell}$. Since all $x_{i}$ are neighbours of $v$, the pseudo-reduction $(C,D)$ induced by $\ell$ is given by $C:=B_{\ell}+v$ and $D:=\{x_{1},x_{2},x_{3},v\}$.
@@
It is straightforward to check that each $(C,D-y)$ with $y\in\{x_{2},x_{3},v\}$ is a tri-separation, and that these cross for different values of $y$. ∎
@@
Assume Section 1.7. A vertex of $O$ is red if it is adjacent to $v$ or incident with two bold edges of $O$.
@@
(Motivation) In what follows, we offer a way to understand the tri-separations interlacing the tri-star of ${\mathcal{A}}$ in Section 1.7 via red mixed 2-separatörs. They have smaller order and hence are easier to analyse.
@@
Assume Section 1.7. If a strong tri-separation $(C,D)$ of $G$ almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$, then $S(C,D)\cap O$ is a red mixed 2-separatör.
@@
Since $(C,D)$ almost interlaces $\sigma$, the elements of the separator $S(C,D)$ are vertices or edges of $O+v$ or edges incident with $v$. By Section 2.4, $S(C,D)$ contains $v$ or an edge incident with $v$. By Section 1.2, no two edges in $S(C,D)$ share ends. Thus, $X:=S(C,D)\cap O$ has size two.
Claim 2.5.0.2.
@@
Proof of Claim. Suppose not for a contradiction. Then one of the intervals of $O_{X}$ is equal to the set of interior points of a timid edge $e=xy$. So one of the sides of $(C,D)$, say $D$, contains all vertices of $O$. So $C$ intersects $V(O)$ precisely in the vertices $x$ and $y$. As the leaf-bags of ${\mathcal{A}}$ are 2-connected and $e$ is timid, the graph $G-v-x-y$ is connected; hence $C$ contains no other vertex of $G-v$. Since $C\setminus D$ is nonempty, it must contain a vertex and the only possibility is that $C\setminus D=\{v\}$. By Section 1.2 the vertex $v$ has at most one neighbour in $D\setminus C=D-x-y$. So by Section 1.7, at most one edge of $O$ can be bold. Thus one of the vertices $x$ and $y$, say $x$, is incident with two timid edges. So $x$ has degree at most three. As $x\in C\cap D$, this contradicts the assumption that $(C,D)$ is strong. $\diamondsuit$
By Claim 2.5.0.2, $X$ is a mixed 2-separatör. It remains to show that $X$ is red. Every edge in $X$ is timid, so let $x$ be a vertex in $X$. Since we are done otherwise, assume that $x$ is not adjacent to $v$. If $x$ is not adjacent to a vertex $y$ in $X$, then the fact that it has at least two neighbours in the sides $C$ and $D$ implies that both its incident edges on $O$ must be bold. So assume that $x$ has a neighbour $y$ in $X$; that is, $X=\{x,y\}$. The only way this is possible is that $xy=:e$ is an edge of $O$. By Claim 2.5.0.2, the edge $e$ must be bold. Let $f$ be the edge of $O$ incident with $x$ aside from $e$.
@@
Assume Section 1.7. For every red mixed 2-separatör $X$ of $G$, there is a tri-separation $(C,D)$ of $G$ that almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$ and satisfies $S(C,D)\cap O=X$.
@@
Denote the intervals of $O_{X}$ by $C_{1}$ and $D_{1}$. We obtain $C_{2}$ from $C_{1}$ by replacing every bold edge $e$ of $O$ with interior in $C_{1}$ by the leaf-side $A_{i}$ of the tri-separation $(A_{i},B_{i})\in\sigma$ that corresponds to $e$, and adding $v$. We define $D_{2}$ analogously. Since $X$ is red, it contains no bold edges and $C_{2}\setminus D_{2}$ and $D_{2}\setminus C_{2}$ are nonempty. Thus $(C_{2},D_{2})$ is a mixed-separation of $G$ that almost interlaces $\sigma$. Its separator is $X+v$, so it is a mixed 3-separation. Vertices of $X$ are red, so have two neighbours in $C_{2}$ and $D_{2}$. By 3-connectivity, $v$ has a neighbour in $C_{2}\setminus D_{2}$ and in $D_{2}\setminus C_{2}$. So if $v$ has a neighbour in $X$, the mixed 3-separation $(C_{2},D_{2})$ is the desired tri-separation. Otherwise the reduction $(C,D)$ of $(C_{2},D_{2})$ satisfies $S(C,D)=X$ and almost interlaces $\sigma$ by Section 2.4, so it is the desired tri-separation. ∎
@@
Assume Section 1.7 with crossing tri-separations and that $O$ does not have the type btt. If a tri-separation $(C,D)$ of $G$ almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$ and $S(C,D)\cap O$ is equal to the boundary of a bold edge of $O$, then $(C,D)$ or $(D,C)$ is in the tri-star of ${\mathcal{A}}$.
@@
Let $e$ be a bold edge of $O$ such that $S(C,D)\cap O$ is equal to the boundary of $e$; if there are two choices for $e$, we denote the other choice by $f$. Let $(A,B)$ be the tri-separation in $\sigma$ that corresponds to the edge $e$. As $(C,D)$ almost interlaces $\sigma$, we have $(A,B)\leqslant(C,D)$ or $(A,B)\leqslant(D,C)$, say the former. Abbreviate $X:=S(C,D)\cap O$. Let $P$ denote the interval of $O_{X}$ that does not include the interior of $e$.
Claim 2.5.0.3.
Possibly after exchanging the roles of ‘$e$’ and ‘$f$’ and adjusting $P$, the sum of neighbours of $v$ on $P$ plus bold edges of $P$ is at least two.
Proof of Claim. Suppose first for a contradiction that $P$ contains no bold edges and at most one neighbour of $v$. Then all edges of $O$ except for $e$ are timid. So all vertices not incident with $e$ are neighbours of $v$ by Section 2.5 and they are in $P$. So $O$ has length three and the type btt. By assumption this type is excluded, so we reach a contradiction.
@@
It remains to suppose for a contradiction that $P$ contains no neighbour of $v$ and exactly one bold edge. If a vertex of $O$ is incident with two timid edges, this vertex is a neighbour of $v$ by Section 2.5 and is in $P$, which is excluded. So every vertex of $O$ is incident with a bold edge. So $O$ is a cycle with exactly two bold edges such that all its vertices are incident with a bold edge. So $O$ contains at most four vertices. By Section 1.7 and since $O$ does not have the type btt by assumption, $O$ has the type ${\tt btbt}^{+}$. As $P$ contains no neighbour of $v$, by Section 1.7 the two endvertices of $e$ or of $f$ are adjacent to $v$. Hence, after possibly exchanging the roles of ‘$e$’ and ‘$f$’ and adjusting $P$, we find that $P$ contains two neighbours of $v$. $\diamondsuit$
@@
By Claim 2.5.0.3 we may assume that the sum of neighbours of $v$ on $P$ plus bold edges of $P$ is at least two. So $v$ has two neighbours in $D$ by Section 1.7. By Section 1.2, two edges incident with $v$ cannot both be in $S(C,D)$, so $v\in D$.
@@
We shall show that $(A,B)=(C,D)$. When restricting to $G-v$, this equality is immediate. By definition of $\sigma$, the vertex $v$ is in $B$. So $B=D$. Since $(A,B)\leqslant(C,D)$, it remains to show that if $v\in C$, then $v\in A$. So assume $v\in C$. Since $(C,D)$ is a tri-separation, $v$ has two neighbours in $C$. As $C-v=A-v$, the vertex $v$ has two neighbours in $A-v$. So by the definition of $(A,B)$, the vertex $v$ is in $A$. This completes the proof. ∎
@@
We say that a mixed 2-separatör $X$ is crossed by a mixed 2-separatör $Y$ if the two intervals of $O_{X}$ contain elements of $Y$; note that crossing is a symmetric relation for mixed 2-separatörs.
@@
Assume Section 1.7. A mixed 2-separatör $X$ of $O$ is crossed by a red mixed 2-separatör of $O$ if and only if $X$ is not equal to the boundary of a bold edge.
@@
Assume Section 1.7 with crossing tri-separations. The tri-star of ${\mathcal{A}}$ is a splitting star of the set of totally-nested nontrivial tri-separations.
@@
Let $(C,D)$ be a nontrivial totally nested tri-separation of $G$ that almost interlaces the tri-star of ${\mathcal{A}}$. Since every tri-separation with a degree-3-vertex $x$ in its separator is crossed (by the atomic cut at $x$), the totally nested tri-separation $(C,D)$ is strong. By Section 2.5, $X:=S(C,D)\cap O$ is a red mixed 2-separatör. By Section 2.5, either $X$ is crossed by a red mixed 2-separator $Y$ or $X$ is equal to the boundary of a bold edge of $O$. In the second case, by Section 2.5 and since we are otherwise done by Section 2.5, the tri-separation $(C,D)$ or $(D,C)$ is in the tri-star of ${\mathcal{A}}$. In the first case, by Section 2.5 there is a tri-separation $(E,F)$ almost interlacing the tri-star of ${\mathcal{A}}$ such that $S(E,F)\cap O=Y$. Since $X$ and $Y$ cross on $O$, the sets $X$ and $Y$ together ensure that all four links of the corner diagram for the tri-separations $(C,D)$ and $(E,F)$ are nonempty; thus the tri-separations $(C,D)$ and $(E,F)$ cross. We have shown that any nontrivial tri-separation $(C,D)$ that interlaces the tri-star of ${\mathcal{A}}$ is crossed by a tri-separation. To summarise, the tri-star of ${\mathcal{A}}$, which consists of totally-nested nontrivial tri-separations by Section 1.7, is splitting within the set of all totally-nested nontrivial tri-separations by Section 2.2.2. ∎
@@
An edge $xy$ of $O$ is good if there is some $(A,B)\in\sigma$ such that $\{x,y\}\subseteq A\cap B$ and $v$ is not in the leaf-side $A$. Let $X$ denote the expanded torso of $\sigma$. The graph $X-v$ is isomorphic to the graph obtained from $O$ by adding for every good edge $xy$ of $O$ a new vertex and joining it to $x$ and $y$ (so that the three vertices form a triangle). Hence $X$ is a concrete generalised wheel with rim $O$. By Section 2.2.1, $X$ is also a generalised wheel, with the same rim. The compressed torso of $\sigma$ is obtained from $X$ by contracting all edges that join $v$ to newly added vertices. Since $X$ is a concrete generalised wheel with rim $O$, it follows that the compressed torso is a wheel with rim $O$. ∎
Proof of Theorem 2.2.1 (ii).
Let $G$ be a 3-connected graph, $N$ its set of totally-nested nontrivial tri-separations, and $\sigma$ a splitting star of $N$. Suppose that $\sigma$ is heavily interlaced by a tri-separation $(A,B)$ of $G$. So $(A,B)$ is half-connected, strong and nontrivial. As $(A,B)$ is not in $N$, it is crossed by a tri-separation $(C,D)$ of $G$. By Section 1.3, we may assume that $(C,D)$ is nontrivial and strong. By the Nested or crossed: analysing corner diagrams (1.3), $(A,B)$ and $(C,D)$ cross so that the centre consists of a single vertex $v$ and all links have size one.
By Section 1.7, $G$ has a 2-connected apex-decomposition ${\mathcal{A}}$ with centre $v$ whose tri-star $\sigma^{\prime}$ is interlaced by $(A,B)$ and $(C,D)$, and whose central torso-cycle $O$ alternates between $S(A,B)-v$ and $S(C,D)-v$; that is to say that we may assume Section 1.7 with crossing tri-separations. By Section 2.5, the tri-star $\sigma^{\prime}$ is a splitting star of $N$. By Section 2.2.2, the fact that $(A,B)$ interlaces the two splitting stars $\sigma$ and $\sigma^{\prime}$ implies $\sigma^{\prime}=\sigma$. Finally, by Section 2.5, the compressed and expanded torsos of $\sigma^{\prime}=\sigma$ are a wheel and generalised wheel, respectively. ∎
2.6. Proof of (iii)

### v3

@@ 4 Tools to prove (ii) and (iii)
We say that a mixed-separation+ $(C,D)$ of a graph $G$ almost interlaces a star $\sigma$ of mixed-separations+ of $G$ if $(A,B)\leqslant(C,D)$ or $(A,B)\leqslant(D,C)$ for all $(A,B)\in\sigma$.
The notion of ‘almost interlaces’ is more general than the notion of ‘interlaces’ in two ways: on the one hand, we consider mixed-separations+, and on the other hand we no longer require that $(A,B)$ and its inverse are not in $\sigma$.
@@
Assume Section 8. If a tri-separation $(C,D)$ of $G$ almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$, then $S(C,D)$ contains $v$ or an edge incident with $v$.
@@
Similar to Section 4, we prove the following (differences are underlined).
@@
5 Proof of (ii)
A key step in the proof of (ii) is to show that the tri-star of the apex-decomposition from Section 8 is splitting, see Section 5 below. Then we finish the proof similarly to the proof of (i). We start preparing to prove Section 5.
@@ 5 Proof of (ii)
Assume Section 8 with crossing tri-separations. If $O$ has type btt, then the tri-star of ${\mathcal{A}}$ is a splitting star of the set of all totally-nested nontrivial tri-separations of $G$.
@@
By Section 9, the tri-star of ${\mathcal{A}}$ consists of totally-nested nontrivial tri-separations, and $v$ is adjacent to all three vertices of $O$. It remains to show that the tri-star of ${\mathcal{A}}$ is splitting. Let $x_{1},x_{2},x_{3}$ denote the vertices of $O$ so that the edge $x_{2}x_{3}$ is bold. Let $B_{\ell}$ denote the unique leaf-bag of ${\mathcal{A}}$; so $\{x_{2},x_{3}\}$ is the adhesion set of $B_{\ell}$. Since all $x_{i}$ are neighbours of $v$, the pseudo-reduction $(C,D)$ induced by $\ell$ is given by $C:=B_{\ell}+v$ and $D:=\{x_{1},x_{2},x_{3},v\}$.
@@
It is straightforward to check that $(C,D-x_{i})$ is a tri-separation for both $i\in\{2,3\}$, and that $(U,W)=(C,D-y)$ is crossed by $(C,D-x_{i})$ for every $i\in\{2,3\}$ with $x_{i}\neq y$. ∎
@@
Assume Section 8. A vertex of $O$ is red if it is adjacent to $v$ or incident with two bold edges of $O$.
@@
(Motivation) In what follows, we offer a way to understand the tri-separations interlacing the tri-star of ${\mathcal{A}}$ in Section 8 via red mixed 2-separatörs. They have smaller order and hence are easier to analyse.
@@
Assume Section 8. If a strong tri-separation $(C,D)$ of $G$ almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$, then $S(C,D)\cap O$ is a red mixed 2-separatör.
@@
Since $(C,D)$ almost interlaces $\sigma$, the elements of the separator $S(C,D)$ are vertices or edges of $O+v$ or edges incident with $v$. By Section 4, $S(C,D)$ contains $v$ or an edge incident with $v$. By Section 3, no two edges in $S(C,D)$ share ends. Thus, $X:=S(C,D)\cap O$ has size two.
Claim 5.0.2.
@@
Proof of Claim. Suppose not for a contradiction. Then one of the intervals of $O_{X}$ is equal to the set of interior points of a timid edge $e=xy$. So one of the sides of $(C,D)$, say $D$, contains all vertices of $O$. So $C$ intersects $V(O)$ precisely in the vertices $x$ and $y$. As the leaf-bags of ${\mathcal{A}}$ are 2-connected and $e$ is timid, the graph $G-v-x-y$ is connected; hence $C$ contains no other vertex of $G-v$. Since $C\setminus D$ is nonempty, it must contain a vertex and the only possibility is that $C\setminus D=\{v\}$. By Section 3 the vertex $v$ has at most one neighbour in $D\setminus C=D-x-y$. So by Section 8, at most one edge of $O$ can be bold. Thus one of the vertices $x$ and $y$, say $x$, is incident with two timid edges. So $x$ has degree at most three. As $x\in C\cap D$, this contradicts the assumption that $(C,D)$ is strong. $\diamondsuit$
By Claim 5.0.2, $X$ is a mixed 2-separatör. It remains to show that $X$ is red. Every edge in $X$ is timid, so let $x$ be a vertex in $X$. Since we are done otherwise, assume that $x$ is not adjacent to $v$. If $x$ is not adjacent to a vertex $y$ in $X$, then the fact that it has at least two neighbours in the sides $C$ and $D$ implies that both its incident edges on $O$ must be bold. So assume that $x$ has a neighbour $y$ in $X$; that is, $X=\{x,y\}$. The only way this is possible is that $xy=:e$ is an edge of $O$. By Claim 5.0.2, the edge $e$ must be bold. Let $f$ be the edge of $O$ incident with $x$ aside from $e$.
@@
Assume Section 8. For every red mixed 2-separatör $X$ of $G$, there is a tri-separation $(C,D)$ of $G$ that almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$ and satisfies $S(C,D)\cap O=X$.
@@
Denote the intervals of $O_{X}$ by $C_{1}$ and $D_{1}$. We obtain $C_{2}$ from $C_{1}$ by replacing every bold edge $e$ of $O$ with interior in $C_{1}$ by the leaf-side $A_{i}$ of the tri-separation $(A_{i},B_{i})\in\sigma$ that corresponds to $e$, and adding $v$. We define $D_{2}$ analogously. Since $X$ is red, it contains no bold edges and $C_{2}\setminus D_{2}$ and $D_{2}\setminus C_{2}$ are nonempty. Thus $(C_{2},D_{2})$ is a mixed-separation of $G$ that almost interlaces $\sigma$. Its separator is $X+v$, so it is a mixed 3-separation. Vertices of $X$ are red, so have two neighbours in $C_{2}$ and $D_{2}$. By 3-connectivity, $v$ has a neighbour in $C_{2}\setminus D_{2}$ and in $D_{2}\setminus C_{2}$. So if $v$ has a neighbour in $X$, the mixed 3-separation $(C_{2},D_{2})$ is the desired tri-separation. Otherwise the reduction $(C,D)$ of $(C_{2},D_{2})$ satisfies $S(C,D)=X$ and almost interlaces $\sigma$ by Section 4, so it is the desired tri-separation. ∎
@@
Assume Section 8 with crossing tri-separations and that $O$ does not have the type btt. If a tri-separation $(C,D)$ of $G$ almost interlaces the tri-star $\sigma$ of ${\mathcal{A}}$ and $S(C,D)\cap O$ is equal to the boundary of a bold edge of $O$, then $(C,D)$ or $(D,C)$ is in the tri-star of ${\mathcal{A}}$ or $(C,D)$ is crossed by some tri-separation of $G$.
@@
Let $e$ be a bold edge of $O$ such that $S(C,D)\cap O$ is equal to the boundary of $e$. Let $(A,B)$ be the tri-separation in $\sigma$ that corresponds to the edge $e$. As $(C,D)$ almost interlaces $\sigma$, we have $(A,B)\leqslant(C,D)$ or $(A,B)\leqslant(D,C)$, say the former. Abbreviate $X:=S(C,D)\cap O$. Let $P$ denote the interval of $O_{X}$ that does not include the interior of $e$.
Claim 5.0.3.
Assume that $P$ contains no neighbour of $v$ and exactly one bold edge $f$. Then $(A,B)=(C,D)$ or $(C,D)$ is crossed by some tri-separation of $G$.
@@
Proof of Claim. If a vertex of $O$ is incident with two timid edges, this vertex is a neighbour of $v$ by Section 5 and is in $P$, which is excluded. So every vertex of $O$ is incident with a bold edge. So $O$ is a cycle with exactly two bold edges $e,f$ such that all its vertices are incident with a bold edge. So $O$ contains at most four vertices. By Section 8 and since $O$ does not have the type btt by assumption, $O$ has the type ${\tt btbt}^{+}$. As $P$ contains no neighbour of $v$, by Section 8 the two endvertices of $e$ are adjacent to $v$ while no endvertex of $f$ is adjacent to $v$. Then both endvertices of $e$ are red, so $X$ equals the set of endvertices of $e$. Let $B_{e}$ denote the leaf-bag at the bold edge $e$, and let $B_{f}$ denote the leaf-bag at the bold edge $f$. Since both endvertices of $e$ are adjacent to $v$, by the definition of pseudo-reduction we have $A=B_{e}\cup\{v\}$, $B=B_{f}\cup\{v\}\cup X$ and $S(A,B)=X\cup\{v\}$. We clearly have $A-v=C-v$ and $B-v=D-v$. We have seen above that $(A,B)\leqslant(C,D)$. So if $v\in D$, then $(A,B)=(C,D)$.
Now assume that $v\notin D$. Then $C=B_{e}\cup\{v\}$ and $D=B_{f}\cup X$. Hence $S(C,D)=X\cup\{vu\}$, where $u$ is the unique neighbour of $v$ in $B_{f}\setminus V(O)$ (using Section 8). We now find a tri-separation of $G$ that crosses $(C,D)$, as follows. Since the two vertices in $X$ must have at least two neighbours in $D$, the edge $e$ of $O$ must exist in $G$. So $X\cup\{v\}$ spans a triangle in $G$. Let $xy$ and $x^{\prime}y^{\prime}$ denote the two timid edges of $O$ so that $X=\{x,x^{\prime}\}$. Let $U:=B_{e}\cup\{v\}$ and $W:=B_{f}\cup\{v\}\cup\{x^{\prime}\}$. Then $S(U,W)=\{xy,v,x^{\prime}\}$. Since $X\cup\{v\}$ spans an edge, $vx^{\prime}\in E(G)$, so $(U,W)$ is a tri-separation. In the corner-diagram for $(C,D)$ and $(U,W)$, $x$ and $vu$ lie in opposite links, so $(C,D)$ and $(U,W)$ cross as desired. $\diamondsuit$
Claim 5.0.4.
The sum of neighbours of $v$ on $P$ plus bold edges of $P$ is at least two, or $(A,B)=(C,D)$ or $(C,D)$ is crossed by some tri-separation of $G$.
Proof of Claim. Assume not for a contradiction. Then by Claim 5.0.3, $P$ contains no bold edges and at most one neighbour of $v$. Then all edges of $O$ except for $e$ are timid. So all vertices not incident with $e$ are neighbours of $v$ by Section 5 and they are in $P$. So $O$ has length three and the type btt. By assumption this type is excluded, so we reach a contradiction. $\diamondsuit$
@@
By Claim 5.0.4 we may assume that the sum of neighbours of $v$ on $P$ plus bold edges of $P$ is at least two. So $v$ has two neighbours in $D$ by Section 8. By Section 3, two edges incident with $v$ cannot both be in $S(C,D)$, so $v\in D$.
@@
We shall show that $(A,B)=(C,D)$. We immediately get $A-v=C-v$ and $B-v=D-v$. By definition of $\sigma$, the vertex $v$ is in $B$. So $B=D$. Since $(A,B)\leqslant(C,D)$, it remains to show that if $v\in C$, then $v\in A$. So assume $v\in C$. Since $(C,D)$ is a tri-separation, $v$ has two neighbours in $C$. As $C-v=A-v$, the vertex $v$ has two neighbours in $A-v$. So by the definition of $(A,B)$, the vertex $v$ is in $A$. This completes the proof. ∎
@@
We say that a mixed 2-separatör $X$ is crossed by a mixed 2-separatör $Y$ if the two intervals of $O_{X}$ contain elements of $Y$ (these elements can be vertices from $Y$ or edges from $Y$ viewed as open intervals). Note that crossing is a symmetric relation for mixed 2-separatörs.
@@
Assume Section 8. A mixed 2-separatör $X$ of $O$ is crossed by a red mixed 2-separatör of $O$ if and only if $X$ is not equal to the boundary of a bold edge.
@@
Assume Section 8 with crossing tri-separations. The tri-star of ${\mathcal{A}}$ is a splitting star of the set of totally-nested nontrivial tri-separations.
@@
Let $(C,D)$ be a nontrivial totally nested tri-separation of $G$ that almost interlaces the tri-star of ${\mathcal{A}}$. Since every tri-separation with a degree-3-vertex $x$ in its separator is crossed (by the atomic cut at $x$), the totally nested tri-separation $(C,D)$ is strong. By Section 5, $X:=S(C,D)\cap O$ is a red mixed 2-separatör. By Section 5, either $X$ is crossed by a red mixed 2-separator $Y$ or $X$ is equal to the boundary of a bold edge of $O$. In the second case, by Section 5 and since we are otherwise done by Section 5, the tri-separation $(C,D)$ or $(D,C)$ is in the tri-star of ${\mathcal{A}}$. In the first case, by Section 5 there is a tri-separation $(E,F)$ almost interlacing the tri-star of ${\mathcal{A}}$ such that $S(E,F)\cap O=Y$. Since $X$ and $Y$ cross on $O$, the sets $X$ and $Y$ together ensure that all four links of the corner diagram for the tri-separations $(C,D)$ and $(E,F)$ are nonempty; thus the tri-separations $(C,D)$ and $(E,F)$ cross. We have shown that any nontrivial tri-separation $(C,D)$ that interlaces the tri-star of ${\mathcal{A}}$ is crossed by a tri-separation. To summarise, the tri-star of ${\mathcal{A}}$, which consists of totally-nested nontrivial tri-separations by Section 8, is splitting within the set of all totally-nested nontrivial tri-separations by Section 2. ∎
@@
An edge $xy$ of $O$ is good if there is some $(A,B)\in\sigma$ such that $\{x,y\}\subseteq A\cap B$ and $v$ is not in the leaf-side $A$. Let $X$ denote the expanded torso of $\sigma$. The graph $X-v$ is isomorphic to the graph obtained from $O$ by adding for every good edge $xy$ of $O$ a new vertex and joining it to $x$ and $y$ (so that the three vertices form a triangle). Hence $X$ is a concrete generalised wheel with rim $O$. By Section 1, $X$ is also a generalised wheel, with the same rim. The compressed torso of $\sigma$ is obtained from $X$ by contracting all edges that join $v$ to newly added vertices. Since $X$ is a concrete generalised wheel with rim $O$, it follows that the compressed torso is a wheel with rim $O$. ∎
Proof of Theorem 2.1 (ii).
Let $G$ be a 3-connected graph, $N$ its set of totally-nested nontrivial tri-separations, and $\sigma$ a splitting star of $N$. Suppose that $\sigma$ is heavily interlaced by a tri-separation $(A,B)$ of $G$. So $(A,B)$ is half-connected, strong and nontrivial. As $(A,B)$ is not in $N$, it is crossed by a tri-separation $(C,D)$ of $G$. By Section 4, we may assume that $(C,D)$ is nontrivial and strong. By the Nested or crossed: analysing corner diagrams (4), $(A,B)$ and $(C,D)$ cross so that the centre consists of a single vertex $v$ and all links have size one.
By Section 8, $G$ has a 2-connected apex-decomposition ${\mathcal{A}}$ with centre $v$ whose tri-star $\sigma^{\prime}$ is interlaced by $(A,B)$ and $(C,D)$, and whose central torso-cycle $O$ alternates between $S(A,B)-v$ and $S(C,D)-v$; that is to say that we may assume Section 8 with crossing tri-separations. By Section 5, the tri-star $\sigma^{\prime}$ is a splitting star of $N$. By Section 2, the fact that $(A,B)$ interlaces the two splitting stars $\sigma$ and $\sigma^{\prime}$ implies $\sigma^{\prime}=\sigma$. Finally, by Section 5, the compressed and expanded torsos of $\sigma^{\prime}=\sigma$ are a wheel and generalised wheel, respectively. ∎
6 Proof of (iii)